OpenStudy (anonymous):

Find the partial derivatives of the function w=Sqrt(4r^2+4s^2+5t^2) ∂w∂r= ∂w∂s= ∂w∂t=

OpenStudy (zarkon):

\[\frac{\partial w}{\partial r}=\frac{8r}{2\sqrt{4r^2+4s^2+5t^2}}\]

OpenStudy (zarkon):

the others are almost Identical